# Skew braces with no proper left ideals

### v1

@@ 1. Introduction
Let $A=(A,\cdot,\circ)$ be a finite skew brace such that $(A,\cdot)\simeq\mathbb{F}_{p}^{n}$ for some prime $p$ and $n\in\mathbb{N}$. Then $A$ is left-simple if and only if $n=1$.
@@
as a linear representation of the finite $p$-group $(A,\circ)$ over the field $\mathbb{F}_{p}$ that has characteristic $p$. The left ideals of the skew brace $A$ then coincide with the $\lambda$-invariant subspaces of the vector space $(A,\cdot)\simeq\mathbb{F}_{p}^{n}$. This means that $A$ is left-simple if and only if $\lambda$ is irreducible. But it is well-known in modular representation theory (e.g. see [11, Proposition 6.2.1]) that $\lambda$ is irreducible if and only if it is the trivial representation. This proves the theorem. ∎
@@
Up to isomorphism, the finite left-simple skew braces with abelian additive group are exactly $(\mathbb{F}_{p},+,+)$, where $p$ is any prime and ++ is the usual addition.
@@ 2. Preliminaries
of their images. The next proposition is essentially [10, Proposition 2.1] and it implies that $\Gamma$ is subgroup of $\mathrm{Aut}(A,\cdot)$ by the normality of $\mathrm{Inn}(A,\cdot)$.
@@
that we can construct. They are implicit in [1, Proof of Theorem A].
@@
Clearly $J_{1}$ is a subgroup of $(A,\cdot)$. For any $a\in A$ and $j\in J_{1}$, we may write $\mathrm{conj}(j)=\lambda_{b}$ for some $b\in A$. It follows that
@@
and so $J_{2}$ is also a subgroup of $(A,\cdot)$. The above further mplies that
@@ 3.1. The case when $n=1$
Conversely, suppose that $A$ is left-simple. Then $A$ is non-trivial, namely, $a\circ b\neq a\cdot b$ for some $a,b\in A$, for otherwise any subgroup of $(A,\cdot)$ would be a left ideal of $A$. We suppose for contradiction that $A$ is not almost trivial. Since $A$ is neither trivial nor almost trivial, we deduce from [3] that $(A,\circ)$ is not simple (the result of [3] is stated in terms of Hopf–Galois structures, but one can easily translate it into skew braces, using [9], for example). We then have $(A,\circ)\not\simeq(A,\cdot)$. By order consideration, it follows that
@@ 2. $\bullet$
Fortunately, the core-free factorizations of finite almost simple groups with a solvable factor have been classified in [8, Theorem 1.1]. Put $H:=\mathrm{Im}(\lambda)$, which is solvable, and $K:=\mathrm{Im}(\rho)$. Identifying $(A,\cdot)$ with $T$, we have:
@@
We now use [8] to show that all possibilities of $(H,K)$ are impossible.
Suppose first that $K$ is also solvable. By [8, Proposition 4.1], for
@@
We have shown that $K$ must be insolvable. From here, we argue depending on the type of $T$. Let us remark that $T$ cannot be an exceptional group of Lie type by [8, Proposition 4.2].
@@ III. $T$ is a classical group of Lie type and is not an alternating group
The possibilities for $(H,K)$ are given in [8, Tables 1.1 and 1.2]. However, they only give an upper bound for $H\cap\mathrm{Inn}(T)$, which is not enough for us. For [8, Table 1.1], we shall further apply [2, Theorem 1], which provides a lower bound for $|H|$ and we can derive a contradiction to (2). For [8, Table 1.2], we shall use the description of $K\cap\mathrm{Inn}(T)$ and obtain a contradiction to (4). We summarize all of the possibilities in the two tables below.
@@
In the first table, write $q=p^{f}$, where $p$ is a prime. As listed in [8, Table 2.1], for example, we know that:
@@
which again contradicts (2). If $n\geq 3$ with $(fn,p)\neq(6,2)$, then we deduce from Zsigmondy’s theorem that $p^{fn}-1$ has a primitive prime divisor $\ell$. In other words, the prime $\ell$ is such that
@@
for all $a\in A$ by (2.1). Observe that
$\mathrm{Im}(\lambda)\not\subseteq\mathrm{Inn}(A,\cdot),\mbox{ or equivalently }\mathrm{Im}(\rho)\not\subseteq\mathrm{Inn}(A,\cdot),$
for otherwise $T\times 1\times\cdots\times 1$ would be a proper left ideal of $A$. We see that $J_{1}\neq A$ by order consideration, and $J_{2}\neq A$. But $J_{1},J_{2}$ are left ideals of $A$ by Proposition 2.2, so the left-simplicity of $A$ implies that $J_{1}=J_{2}=1$.
  1. $\bullet$
@@
  2. $\bullet$
    Since $J_{2}=1$, the circle group $(A,\circ)$ is isomorphic to the subgroup $\mathrm{Im}(\rho)$ of $\mathrm{Aut}(A,\cdot)\simeq\mathrm{Aut}(T)^{n}\rtimes S_{n}$. The projection $P$ of $\mathrm{Im}(\rho)$ onto $S_{n}$ is equal to that of $\mathrm{Im}(\lambda)$ by (3.4). Thus, for any orbit $O$ of the natural action of $P$ on $\{1,2,\dots,n\}$, the subgroup of $(A,\cdot)$ defined by
@@ 2. $\bullet$
    where $T_{i}$ denotes the $i$th copy of $T$ in $(A,\cdot)\simeq T^{n}$, is a left ideal of $A$. But $A$ is left-simple, so necessarily $O=\{1,2,\dots,n\}$. This proves that $P$ is a transitive subgroup of $S_{n}$.
@@ 4. Connection with minimal Hopf–Galois structures
$L^{\mathcal{H}^{\prime}}=\{x\in L\mid h^{\prime}(x)=\epsilon(h^{\prime})x\mbox{ for all }h^{\prime}\in\mathcal{H}^{\prime}\}$
@@
that is analogous to the Galois correspondence
@@
$H$ $\longmapsto L^{H}$
in Galois theory. Indeed, when $L/K$ is a Galois extension with Galois group $G$, the group ring $K[G]$ is naturally a Hopf–Galois structure on $L/K$, and the map $\Phi_{K[G]}$ reduces to the above Galois correspondence.
It is known by [4] that the map $\Phi_{\mathcal{H}}$ is always injective, but it need not be surjective in general. We can of course ask when $\Phi_{\mathcal{H}}$ is surjective. Here, we consider the other extreme and ask when $\Phi_{\mathcal{H}}$ is as far from being surjective as possible, i.e. when $\mathrm{Im}(\Phi_{\mathcal{H}})$ only contains $L$ and $K$. Since $\Phi_{\mathcal{H}}$ is injective, this is precisely the situation when $K$ and $\mathcal{H}$ are the only $K$-subalgebras of $\mathcal{H}$. The next definition is due to [6].
@@
The theory of Hopf–Galois structures applies to all extensions. But when restricted to separable extensions, we have a very nice group-theoretic classification of Hopf–Galois structures thanks to [7]. In addition, when further restricted to Galois extensions, the classification may be translated into the language of skew braces. The next two theorems are due to [9].
@@ 2. $(2)$
$L[(G,\cdot)]^{(G,\circ)}=\left\{\sum_{\tau\in G}\ell_{\tau}\tau:L[(G,\cdot)]:\sigma(\ell_{\tau})=\ell_{\lambda_{\sigma}(\tau)}\mbox{ for all }\sigma,\tau\in G\right\}$
@@ Acknowledgements
This work is supported by JSPS KAKENHI 24K16891.

### v2

@@ 1. Introduction
Let $A=(A,\cdot,\circ)$ be a finite skew brace such that $(A,\cdot)\simeq C_{p}^{n}$ for some prime $p$ and $n\in\mathbb{N}$. Then $A$ is left-simple if and only if $n=1$.
@@
as a linear representation of the finite $p$-group $(A,\circ)$ over the field $\mathbb{F}_{p}$ that has characteristic $p$. The left ideals of the skew brace $A$ then coincide with the $\lambda$-invariant subspaces of the vector space $(A,\cdot)\simeq\mathbb{F}_{p}^{n}$. This means that $A$ is left-simple if and only if $\lambda$ is irreducible. But it is well-known in modular representation theory (e.g. see [13, Proposition 6.2.1]) that $\lambda$ is irreducible if and only if it is the trivial representation. This proves the theorem. ∎
@@
Up to isomorphism, the finite left-simple skew braces with an abelian additive group are exactly $\mathbb{F}_{p}=(\mathbb{F}_{p},+,+)$, where $p$ is any prime and ++ is the usual addition.
@@ 2. $(b)$
Remark 1.6.
A skew brace $A=(A,\cdot,\circ)$ is said to be simple if $A\neq 1$ and it has no ideal other than $1$ and $A$. Recall that an ideal of $A$ is a left ideal $I$ of $A$ that is normal in both $(A,\cdot)$ and $(A,\circ)$. In this case, we can naturally define a quotient skew brace structure on the set
$A/I=\{a\cdot I:a\in A\}=\{a\circ I:a\in A\}.$
Thus, ideals in a skew brace are a natural analog of normal subgroups in a group, and simple skew braces are arguably a closer analog of simple groups than left-simple skew braces. A classification of finite simple skew braces is desirable, but it seems to be out of reach at the moment. For example, in a simple skew brace $A=(A,\cdot,\circ)$, it is possible that $(A,\cdot)$ is abelian [1].
@@ 2. Preliminaries
of their images. The next proposition is essentially [12, Proposition 2.1] and it implies that $\Gamma$ is subgroup of $\mathrm{Aut}(A,\cdot)$ by the normality of $\mathrm{Inn}(A,\cdot)$.
@@
that we can construct. They are implicit in [2, Proof of Theorem A].
@@
Clearly $J_{1}$ is a subgroup of $(A,\cdot)$. Let $j\in J_{1}$ and write $\mathrm{conj}(j)=\lambda_{b}$ for some $b\in A$. For any $a\in A$, it then follows that
@@
and so $J_{2}$ is also a subgroup of $(A,\cdot)$. The above further implies that
@@ 3.1. The case when $n=1$
Conversely, suppose that $A$ is left-simple. Then $A$ is non-trivial, namely, $a\circ b\neq a\cdot b$ for some $a,b\in A$, for otherwise any subgroup of $(A,\cdot)$ would be a left ideal of $A$. We suppose for contradiction that $A$ is not almost trivial. Since $A$ is neither trivial nor almost trivial, we obtain from [4, Theorem 1.1] that $(A,\circ)$ is not simple (the result of [4] is stated in terms of Hopf–Galois structures, but using [11, Theorem 3.1], for example, to translate it into the language of skew braces, it says that a finite skew brace with a non-abelian simple multiplicative group is either trivial or almost trivial). We then have $(A,\circ)\not\simeq(A,\cdot)$. By order consideration, it follows that
@@
Indeed, note that $\mathrm{Inn}(A,\cdot)\simeq(A,\cdot)$ has order $|A|$, while
$\mathrm{Im}(\lambda)\simeq(A,\circ)/\ker(\lambda)\quad\mbox{and}\quad\mathrm{Im}(\rho)\simeq(A,\circ)/\ker(\rho)$
have order at most $|A|$. This implies that
[…]
and we have a contradiction in both cases.
@@ 2. $\bullet$
Fortunately, the core-free factorizations of finite almost simple groups with a solvable factor have been classified in [9, Theorem 1.1]. Put $H:=\mathrm{Im}(\lambda)$, which is solvable, and $K:=\mathrm{Im}(\rho)$. Identifying $(A,\cdot)$ with $T$, we have:
@@
We now use [9] to show that all candidates of $(H,K)$ are impossible.
Remark 3.1.
In the following tables, the orders of certain groups are listed. For the linear groups, they can be constructed in Magma via
  1. $\bullet$
    PSL(n,q) for the projective special linear group on $\mathrm{GF}(q)^{n}$;
  2. $\bullet$
    PSU(n,q) for the projective special unitary group on $\mathrm{GF}(q^{2})^{n}$;
  3. $\bullet$
    PSp(2n,q) for the projective sympletic group on $\mathrm{GF}(q)^{2n}$;
  4. $\bullet$
    PSO(2n+1,q) for the projective special orthogonal group on $\mathrm{GF}(q)^{2n+1}$;
    (the symbol $\mathrm{P}\Omega_{2n+1}(q)$ denotes its derived subgroup);
  5. $\bullet$
    PSOPlus(2n,q) and PSOMinus(2n,q), respectively, for the projective special orthogonal groups on $\mathrm{GF}(q)^{2n}$ with Witt indices $n$ and n1n-1;
    (the symbols $\mathrm{P}\Omega_{2n}^{+}(q)$ and $\mathrm{P}\Omega_{2n}^{-}(q)$ denote their derived subgroups);
  6. $\bullet$
    AGammaL(n,q) for the affine semilinear group on $\mathrm{GF}(q)^{n}$;
  7. $\bullet$
    ChevalleyGroup("G",2,q) for the Chevalley group or exceptional group of type $G_{2}$ over the field $\mathrm{GF}(q)$.
For the non-projective linear groups, simply remove P in the above. For the Mathieu groups, they can be constructed in GAP via
  1. $\bullet$
    MathieuGroup(n) for the Mathieu group acting on $n$ objects.
The orders of these groups may then be computed. The orders of the outer automorphism groups $\mathrm{Out}(T)$ may be computed via
$|\mathrm{Out}(T)|=|\mathrm{Aut}(T)|/|\mathrm{Inn}(T)|=|\mathrm{Aut}(T)|/|T|.$
For the symmetric and alternating groups, their orders are obvious, so we do not list them explicitly.
Suppose first that $K$ is also solvable. By [9, Proposition 4.1], for
@@ 1. $\bullet$
We have shown that $K$ must be insolvable. From here, we argue depending on the type of $T$. Let us remark that $T$ cannot be an exceptional group of Lie type by [9, Proposition 4.2].
@@ III. $T$ is a classical group of Lie type and is not an alternating group
The possibilities for $(H,K)$ are given in [9, Tables 1.1 and 1.2]. However, they only give an upper bound for $H\cap\mathrm{Inn}(T)$, which is not enough for us. For [9, Table 1.1], we shall further apply [3, Theorem 1], which provides a lower bound for $|H|$ and we can derive a contradiction to (2). For [9, Table 1.2], we shall use the description of $K\cap\mathrm{Inn}(T)$ and obtain a contradiction to (4). We summarize all of the possibilities in the two tables below.
@@
In the first table, write $q=p^{f}$, where $p$ is a prime. As listed in [9, Table 2.1], for example, we know that:
@@
which again contradicts (2). If $n\geq 3$ with $(fn,p)\neq(6,2)$, then we deduce from Zsigmondy’s theorem [14] (also see [10, Theorem 3]) that $p^{fn}-1$ has a primitive prime divisor $\ell$. In other words, the prime $\ell$ is such that
@@
for all $a\in A$ by (2.1), and that
$\mathrm{Im}(\lambda)\not\subseteq\mathrm{Inn}(A,\cdot),$ (3.5)
for otherwise $T\times 1\times\cdots\times 1$ would be a non-trivial proper left ideal of $A$. We then see that $J_{1}\neq A$ by order consideration. Indeed, if $J_{1}=A$, namely the inclusion $\mathrm{Inn}(A,\cdot)\subseteq\mathrm{Im}(\lambda)$ holds, then we must in fact have an equality because $\mathrm{Inn}(A,\cdot)\simeq(A,\cdot)$ has order $|A|$, while
$\mathrm{Im}(\lambda)\simeq(A,\circ)/\ker(\lambda)$
has order at most $|A|$. But this contradicts (3.5). From (3.4) and (3.5), we also obtain $J_{2}\neq A$. Since $J_{1},J_{2}$ are left ideals of $A$ by Proposition 2.2, the left-simplicity of $A$ implies that $J_{1}=J_{2}=1$.
@@
Since $J_{2}=1$, the multiplicative group $(A,\circ)$ embeds into
$\mathrm{Aut}(A,\cdot)\simeq\mathrm{Aut}(T)^{n}\rtimes S_{n}$
via $\rho$, and let $P$ denote the projection of $\mathrm{Im}(\rho)$ onto $S_{n}$. Since
$\mathrm{Inn}(A,\cdot)\simeq\mathrm{Inn}(T)^{n}\subseteq\mathrm{Aut}(T)^{n},$
the projection of $\mathrm{Im}(\lambda)$ onto $S_{n}$ is also equal to $P$ by (3.4). For any orbit $O$ of the natural action of $P$ on $\{1,2,\dots,n\}$, the subgroup of $(A,\cdot)$ given by
@@
where $T_{i}$ denotes the $i$th copy of $T$ in $(A,\cdot)\simeq T^{n}$, is then a left ideal of $A$. But $A$ is left-simple, so necessarily $O=\{1,2,\dots,n\}$. This proves that $P$ is a transitive subgroup of $S_{n}$.
@@
Remark 3.2.
The condition that $(A,\circ)$ is isomorphic to a subgroup, $X$ say, of $\mathrm{Aut}(T^{n})$ whose projection onto $S_{n}$ is transitive is very restrictive. To see why, consider the natural homomorphism
$\phi:X\lhook\joinrel\xrightarrow{\quad}\mathrm{Aut}(T)^{n}\rtimes S_{n}\longrightarrow\mathrm{Out}(T)^{n}\rtimes S_{n}.$
Its kernel $\ker(\phi)$ lies in $\mathrm{Inn}(T)^{n}\simeq T^{n}$, and its image $\mathrm{Im}(\varphi)\simeq X/\ker(\phi)$ has the same projection onto $S_{n}$ as $X$, which is transitive. It follows that
$\exists R\leq T^{n}\mbox{ s.t. }n\mid[T^{n}:R]\mbox{\,\ and\,\ }[T^{n}:R]\mid|\mathrm{Out}(T)|^{n}\cdot n!.$ (3.6)
In particular, the prime factors of $n$ must all divide $|T|$.
For example, take $T=(T,\cdot)=(A_{5},\cdot)$. Note that
$|A_{5}|=60=2^{2}\cdot 3\cdot 5\quad\mbox{and}\quad|\mathrm{Out}(A_{5})|=2.$
We see right away that $(A_{5}^{n},\cdot,\circ)$ can be left-simple only when $n=2^{x}3^{y}5^{z}$, where $x,y,z\in\mathbb{N}_{\geq 0}$. Moreover, a skew brace $(A_{5}^{2},\cdot,\circ)$ cannot be left-simple because $A_{5}^{2}$ has no subgroup of index $2,4$, or $8$.
@@ 4. Connection with minimal Hopf–Galois structures
$L^{\mathcal{H}^{\prime}}=\{x\in L:h^{\prime}(x)=\epsilon(h^{\prime})x\mbox{ for all }h^{\prime}\in\mathcal{H}^{\prime}\}$
@@
that is analogous to the Galois correspondence in Galois theory. Indeed, in the case that $L/K$ is a $G$-Galois extension, the group ring $K[G]$ is naturally a Hopf–Galois structure on $L/K$, and the map $\Phi_{K[G]}$ above reduces to the usual Galois correspondence
@@
$G^{\prime}$ $\longmapsto L^{G^{\prime}}$
because the Hopf $K$-subalgebras of $K[G]$ are exactly the group rings $K[G^{\prime}]$, where $G^{\prime}$ ranges over the subgroups of $G$.
It is known by [5] that the map $\Phi_{\mathcal{H}}$ is always injective, but it need not be surjective in general. We can of course ask when $\Phi_{\mathcal{H}}$ is surjective. Here, we consider the other extreme and ask when $\Phi_{\mathcal{H}}$ is as far from being surjective as possible, i.e. when $\mathrm{Im}(\Phi_{\mathcal{H}})$ only contains $L$ and $K$. Since $\Phi_{\mathcal{H}}$ is injective, this is precisely the case when $K$ and $\mathcal{H}$ are the only Hopf $K$-subalgebras of $\mathcal{H}$. The next definition is due to [7].
@@
The theory of Hopf–Galois structures applies to all extensions. But when restricted to separable extensions, we have a very nice group-theoretic classification of Hopf–Galois structures thanks to [8]. In addition, when further restricted to Galois extensions, the classification may be translated into the
language of skew braces. The next theorem is from [11, Theorem 3.1].
@@ 2. $(2)$
$L[(G,\cdot)]^{(G,\circ)}=\left\{\sum_{\tau\in G}\ell_{\tau}\tau\in L[(G,\cdot)]:\sigma(\ell_{\tau})=\ell_{\lambda_{\sigma}(\tau)}\mbox{ for all }\sigma,\tau\in G\right\}$
@@
Let us remark that the Hopf–Galois structure corresponding to the trivial skew brace is the group ring $K[G]$ and is called the classical structure, while the Hopf–Galois structure corresponding to the almost trivial skew brace is called the canonical non-classical structure.
The next theorem is from [11, Corollary 4.1].
@@ Acknowledgements
This work is supported by JSPS KAKENHI 24K16891. The author would like to thank the referees for their valuable suggestions.